The lengths of the tangents from A(- 3, 5) and B(1, 2) to a circle are 6 and 5 respectively. Then what is the length of the tangent from C(5, - 1) to the same circle?

Grade:11

1 Answers

Saurabh Koranglekar
5 years ago
Dear student

Let the circle centre be ( h,k)

Eq 1. (h +3)^2 +( k-5)^2 = r^2 + 36

Eq 2 ( h-1)^2 + ( k-2)^2 = r^2 + 25

As you can see, 3 variables and 2 equations cant have a unique solution,

Please check and repost the question with correction

Regards